# To Bias or Not to Bias: Detecting bias in News with bias-detector

### v1

@@ 7 Conclusion
In this work, we presented a fine-tuned RoBERTa-based model called bias-detector for sentence-level media bias detection, trained on the expert-annotated BABE dataset. Through comprehensive evaluation and statistical testing, our model outperformed the domain-adapted DA-RoBERTa baseline across multiple folds, achieving higher macro F1-scores and demonstrating statistically significant enhancements. Attention-based analyses further revealed that our model attends more meaningfully to linguistic cues associated with bias.
@@ 8 Acknowledgements
I acknowledge the support and guidance of my Master-degree thesis supervisor Ahmed Mosharafa and NLP Professor Georg Groh, because of whom, this work was possible.

### v2

@@ 7 Conclusion
In this work, we presented a fine-tuned RoBERTa-based model called bias-detector for sentence-level media bias detection, trained on the expert-annotated BABE dataset. Through comprehensive evaluation and statistical testing, our model outperformed the domain-adapted DA-RoBERTa-FT baseline across multiple folds, achieving higher macro F1-score and demonstrating statistically significant enhancements. Attention-based analyses further revealed that our model attends more meaningfully to linguistic cues associated with bias.
@@ 8 Acknowledgements
I acknowledge the support and guidance of my Master-degree thesis supervisor Ahmed Mosharafa and Professor Georg Groh.